Can you please help me with the problem below:
Two points, A and B are 10 units apart. How may points are equidistant from both A and B and also 6 units from Ponit A?
Thanks
Appreciate IT!

Answer by RAY100(1637)

Points A & B are 10 units apart
.
ALL points equidistant, in 2d,,are on the perpendicular bisector of AB
.
If points are 6 units from A,,,,they are also 6 units from B
.
Using the Pythagorean Theorem,c^2 = a^2 +b^2, ,,,on the right triangle
.
legs 5 (10/2) and x,,,and hypotenuse 6,,,
.
6^2 = 5^2 + x^2,,,,36-25 = 11 = x^2
.
x= sqrt 11= +/- 3.317
.
That is, There is a point +3.317 above the midpoint, and another point 3.317 below the midpoint, in 2 Dimension
.
.
In 3 Dimensions, it can be visualized, an infinite numbers of points, in a circle,
.
on the plane perpendicular to the midpoint of AB, with a radius
.
of 3.317
.